# Probabilistic Sampling of Balanced K-Means using Adiabatic Quantum Computing

### v1

@@ 15 Points, 3 Clusters, 2 Dim 10 Points, 2 Clusters, 2 Dim 20 Points, 4 Clusters, 4 Dim
5.3 Calibration Performance
We evaluate the calibration of our method on a synthetic clustering scenario with 3 clusters and 15 points using QA, SIM and exhaustive search on 1000 tasks. First all clustering solutions $Z$ are accumulated in bins according to their estimated posterior probability $P(Z|\mathbf{X})$. This process also includes all sampled non-optimal but feasible solutions. The resulting histogram of solutions is shown in Figure 1(a). After accumulation, the ratio of correct solutions in each bin is evaluated in Figures 1(b),1(c), and 1(d), which for a calibrated method should be close to the mean predicted probability represented by the diagonal. We find our approach to generate well-calibrated probabilities in both simulation and when using QA.
(a) Histogram of predicted probabilities.
(b) 2 clusters, 10 points.
(c) 3 clusters, 15 points.
(d) 4 clusters, 20 points.
Figure 2: Evaluation of the calibration for QA, SIM and exhaustive search in a clustering scenario with 15 points and 3 clusters. Evaluation of the calibration for quantum annealing, simulated annealing and exhaustive search in clustering scenarios with 2, 3, and 4 clusters and 5 points in each cluster. All results are generated with 1,000 problems in each scenario and 5,000 measurements for each clustering problem.
5.4 Clustering Performance
We evaluate the performance of our clustering formulation using synthetic data in Table 1. The upper rows show results for 3 clusters with an increasing number of points and 1000 tasks/setting for SIM and k-means. Due to the problem size, results cannot be found using QA and exhaustive search. Comparing results shows that most clustering metrics are better when using our formulation with SIM, however, the difference decreases with increasing problem size. This can be attributed to the globally optimal solution our approach is optimizing for. While in an ideal the global solution is at least as good as the k-means solution, an increasing problem size also increases the complexity of the corresponding QUBO. This makes it harder to solve exactly and the noisy approach using annealing is not capable of solving the problem correctly.
For the largest clustering problem with 3 clusters and 45 points, k-means provides the best accuracy and thus, a larger number of correct solutions compared to SIM. However, the metrics indicating clustering quality are higher for SIM. This indicates that our formulation is able to find a better solution in cases where the problem is not solved correctly by SIM and k-means.
@@ 5.4 Clustering Performance
The quality of predicted posterior probabilities can be further evaluated by linking them to the clustering metrics and sparsifying the set of tasks. Starting with metrics over the whole set of clustering tasks, we remove tasks according to increasing predicted probability and evaluate the metric over the remaining set. In an ideal predictor, this removes the lowest performing tasks first. In Figures LABEL:fig:sparsification_method_comp and LABEL:fig:sparsification_metric_comp this is done for the adjusted Rand index with different solvers and for QA with different metrics respectively. The solid lines show the sparsification-plots using the predicted probabilities and the dashed lines show the same plots for an oracle method that generates the best possible ordering based on the metrics themselves.
In Figure LABEL:fig:sparsification_method_comp it becomes apparent that QA, SIM and exhaustive search perform close to each other over most of the value range. Nevertheless, for a high sparsification with more than 80% of the tasks removed, QA shows a drop in performance compared to the other methods. This is caused by tasks where only a single, but incorrect solution is found, which gets assigned a posterior probability of $P(Z|\mathbf{X})=1.0$. In such cases the Boltzmann distribution has not been sampled sufficiently well, either because of too few measurements or because of a low effective sampling temperature. As SIM does not show this behavior the source can likely be traced back to current limitation of the quantum computer.
5.5 Maximum Pointsets
Qualitative examples for the maximum pointsets generated with Algorithm 1 are depicted in Figure 4. The shape of each point represents the ground truth class and the color the assigned cluster. Starting with the most likely solution having a predicted probability of $p(Y|X)=0.61$ on the left, each plot shows one additional step of the algorithm, which successively removes points from the solution, indicated by plotting them in Grey. The illustration demonstrates that the MaxsetSearch algorithm is able to generate well-separated clusters from the probabilistic predictions.
@@ 5.5 Maximum Pointsets
Figure 4: Visualization of max pointsets for synthetic data with the probability for each of the determined pointsets.
5.6 IRIS Dataset
While our method assumes an identity covariance in the data, it can be applied to other distributions, which we evaluate on the widely used IRIS dataset. Clustering metrics for experiments using 3 clusters with 5 points each are provided in Table 6 and show that our formulation using QA and SIM is competitive with k-means. Figures LABEL:fig:qual_iris_qa and LABEL:fig:qual_iris_sim show differently sized qualitative examples from the dataset solved using our formulation with QA and SIM respectively.
@@ 5.6 IRIS Dataset
Figure 6: Clustering metrics generated on IRIS subsets using 3 clusters with 15 points in total.
Though the value of the predicted probability cannot be interpreted easily with the mismatch between the assumed and the actual data-generating process, it can be used to find maximum pointsets. The results from Algorithm 1 using results from SIM are provided in Figure 7 and show that even for a distribution mismatch the maximum pointsets can provide meaningful results for removing ambiguous samples.
Figure 7: Max pointsets generated using simulated annealing on the IRIS dataset with 60 points.
@@ 6 Conclusion
In this work, we proposed a probabilistic clustering approach based on sampling k-means solutions using AQC. By using all valid measurements, calibrated confidence scores are computed at little cost and solutions are competitive to an iterative balanced k-means approach. We evaluated our method on synthetic as well as real data using simulation, exhaustive search as well as the D-Wave Advantage 2 prototype AQC to explore the potential of quantum computing in machine learning and computer vision.

### v2

@@ 15 Points, 3 Clusters, 2 Dim 10 Points, 2 Clusters, 2 Dim 20 Points, 4 Clusters, 4 Dim
Data for the quantitative evaluation of our method is synthetically generated. For each clustering problem a total of $I$ points are sampled from a separate normal distribution for each of $K$ clusters. The centroids are randomly drawn, such that the distance between each pair of clusters lies within a predefined range $[d_{\text{min}},d_{\text{max}}]$. For each experiment a total of $L$ clustering tasks is generated. This allows us to evaluate the calibration metrics over a large value range. For all experiments that directly compare methods, identical clustering tasks are used.
Further results are provided for the IRIS dataset [38] and a dataset of images containing ambiguous objects that are to be identified. The IRIS contains 50 samples of 4 features in 3 classes. We randomly subsample the points and dimensions to generate the parameters required for our experiments.
Clustering metrics are computed using the available ground-truth clusters. We evaluate 4 standard metrics: The accuracy, which measures the ratio of clustering solutions that are identical to the ground-truth. Completeness [69] measures the ratio of points from a single cluster being grouped together. The adjusted Rand score [44] compares all pairs of points in the ground truth and prediction and the Fowlkes-Mallows index [40] combines precision and recall into a single score.
5.1 Results
Calibration performance.
We evaluate the calibration of our method on a synthetic clustering scenario with 3 clusters and 15 points using QA, SIM and exhaustive search on 1000 tasks. First all clustering solutions $Z$ are accumulated in bins according to their estimated posterior probability $P(Z|X)$. This process also includes all sampled non-optimal but feasible solutions. Figure 2 shows the resulting histogram of solutions and the ratio of correct solutions in each bin after accumulation. It thus evaluates calibration, where the ideal case has mean predicted probabilities on the diagonal. We find our approach to generate well-calibrated probabilities in both simulation and when using QA for the two smaller experiments for up to 15 points. Due to the problem QUBO size, only SIM is able to find the correct solutions on the largest example.
Clustering performance.
We evaluate the performance of our clustering formulation using synthetic data in Table 1. The upper rows show results for 3 clusters with an increasing number of points and 1000 tasks/setting for SIM and k-means. Due to the problem size, it cannot be solved using QA and exhaustive search. Our formulation with SIM outperforms k-means on the small tasks, however, the difference vanishes with increasing problem size. This can be attributed to the globally optimal solution our approach is optimizing for. Ideally, the global solution is at least as good as the k-means solution, however, an increasing problem size also increases the complexity of the QUBO, which explains the dropping performance on large problems.
For the largest clustering problem with 3 clusters and 45 points, k-means provides the best accuracy with more correct solutions compared to SIM. However, the clustering quality metrics are higher for SIM. This indicates that our formulation is able to find a better solution in cases where the problem is not solved correctly by SIM and k-means.
@@ 5.1 Results
The quality of predicted posterior probabilities can be further evaluated by linking them to the clustering metrics and sparsifying the set of tasks. Starting with metrics over the whole set of clustering tasks, we remove tasks according to increasing probability and evaluate the metric over the remaining set. For an ideal predictor, this removes the lowest-performing tasks first. In Figure LABEL:fig:sparsification_method_comp this is done for the adjusted Rand index with different solvers and in Figure LABEL:fig:sparsification_metric_comp QA with different metrics. The solid lines show the sparsification plots using the predicted probabilities and the dashed lines show the same plots for an oracle method that generates the best possible ordering based on the metrics.
In Figure LABEL:fig:sparsification_method_comp it becomes apparent that QA, SIM and exhaustive search perform close to each other over most of the value range. Nevertheless, for a high sparsification with more than 80% of the tasks removed, QA shows a drop in performance compared to the other methods. This is caused by tasks where only a single, but incorrect solution is found, which gets assigned a posterior probability of $P(Z|X)=1.0$. In such cases, the Boltzmann distribution has not been sampled sufficiently well, either because of too few measurements or because of a low effective sampling temperature. As SIM does not show this behavior the source can likely be traced back to current limitation of the quantum computer.
Coresets.
Qualitative examples for the coresets generated with Algorithm 1 are depicted in Figure 4. The shape of each point represents the ground truth class and the color the assigned cluster. Starting with the most likely solution having a predicted probability of $p(Z|X)=0.61$ on the left, each plot shows one additional step of the algorithm, which successively removes points from the solution, indicated by plotting them in Grey. The illustration demonstrates that the CoresetSearch algorithm is able to generate well-separated clusters from the probabilistic predictions.
@@
Figure 4: Visualization of coresets for synthetic data with the probability for each of the determined pointsets.
{floatrow}
@@
[…]
Figure 6: Performance on IRIS subsets (3 clusters/15 points).
@@
[…]
Real-World Datasets
While our method assumes an identity covariance in the data, it can be applied to other distributions, which we evaluate on the widely used IRIS dataset as well as a set of self-collected images. Clustering metrics for experiments on IRIS using 3 clusters with 5 points each are provided in Table 5(b) and show that our formulation using QA and SIM is competitive with k-means. Figures LABEL:fig:qual_iris_qa and LABEL:fig:qual_iris_sim show differently sized qualitative examples from the dataset solved using our formulation with QA and SIM respectively. On the full IRIS dataset, SIM and k-means achieve identical results with a Completeness of 77.7%, Adjusted Rand index of 78.6% and further results provided in the supplementary. The results from Algorithm 1 using results from SIM are provided in Figures 7 and 8 for IRIS and our dataset respectively. They show that even for a distribution mismatch the coresets can provide meaningful results for removing ambiguous samples.
To demonstrate that meaningful ambiguous samples can be identified in a high-dimensional space present in image data, we collected publicly available images of cars, boats and cars towing boats. Image features were extracted using VGG [72] and subsequently clustered using our approach. Similar to IRIS, the distribution does not match our assumptions, however, we are still able to identify ambiguous images that contain cars towing a boat by using Algorithm 1. The first coreset is depicted in Figure 8 where differently colored frames indicate the cluster assignment and Grey frames mark ambiguous samples. Clustering is performed on the 1000-dimensional feature and images are plotted at their 2D projection generated using UMAP [59]. Our approach correctly identifies the images that contain a car towing a boat and thus, cannot be uniquely assigned to either cluster.
Figure 7: Coresets generated using SIM on the IRIS dataset with 60 points.
Figure 8: First coreset using SIM on our collected dataset.
@@ 6 Conclusion
In this work, we proposed a probabilistic clustering approach based on sampling k-means solutions using AQC. By using all valid measurements, calibrated confidence scores are computed at little cost and solutions are competitive to an iterative balanced k-means approach. We evaluated our method on synthetic as well as real data using simulation, exhaustive search as well as the D-Wave Advantage 2 prototype QA to explore the potential of quantum computing in machine learning and computer vision.